OpenStudy (goformit100):

What is the slope of the line 2x + 5y = -10? negative five halves five halves two fifths negative two fifths none of the above

OpenStudy (goformit100):

@Jonask

OpenStudy (anonymous):

express equation in the form \[y=mx+c\]

OpenStudy (anonymous):

by taking 2x to the right hand side ,then dividing by 5

OpenStudy (anonymous):

ie subtracting 2x both sides

OpenStudy (goformit100):

what's the final answer ?

OpenStudy (goformit100):

@Jonask

OpenStudy (anonymous):

you shud get it, cos we are not here to give out answers but to help you get the answer

OpenStudy (anonymous):

jus follow the instructions

OpenStudy (goformit100):

thanks

OpenStudy (anonymous):

2x + 5y = -10 subtract 2x both sides

OpenStudy (anonymous):

5y=-10-2x y=-10/5-2x/5 y=-2-2x/5 so m=2/5